# Regularization of the superposition principle: Potential theory meets Fokker-Planck equations

### 05 Mar 2026 v1

@@ 1 Introduction
where $W$ is a $d$-dimensional standard Brownian motion whilst $\mathcal{L}_{X(t)}$ denotes the law of $X(t)$, $t\geq 0$. The converse implication, namely that the one-dimensional time marginals of a weak solution to the McKean-Vlasov SDE (1.2) solves the (non)linear Fokker-Planck equation (1.1), is just a simple consequence of Ito’s formula. Thus, under mild assumptions on the coefficients, there is a well-understood correspondence between (1.1) and (1.2). The hard implication is, of course, the one that comes from the superposition principle, which was first obtained in the linear case (i.e. the coefficients do not depend on $\mu_{t}$) by [1] and [46], followed up by [70], [35], as well as by other contributions. The extension of the superposition principle to the nonlinear setting and the corresponding McKean-Vlasov SDEs was realized in [3], [4]. We refer to [36] for a comprehensive study of the theory of linear Fokker-Planck-Kolmogorov equations, to [39] for a detailed presentation of McKean-Vlasov SDEs and their applications, and to [6] for the analysis of nonlinear Fokker-Planck equations and their probabilistic counterpart which are nonlinear Markov processes in the sense of McKean; see [58], [65]. See also [2] for a gradient flow theory in metric spaces and the superposition principle for the continuity equation.
Another fundamental feature of the superposition principle is that, under a weak uniqueness assumption, the martingale solution denoted above by $\eta$ enjoys the ”simple” Markov property; see Section 2.1, more precisely (2.10). Thus, under very mild assumptions on the coefficients, the duality of (non)linear Fokker-Planck equations and McKean Vlasov SDEs could be analyzed through the lens of Markov processes. However, it is well known that the ”simple” Markov property (2.10) is typically not enough in order to benefit from the rich theory of Markov processes, since additional regularity properties of the process, like the strong Markov property, are necessary. This leads us to the formulation of the first main goal which we achieve in this paper:
@@
As indicated above, a right process automatically has the strong Markov property, in particular we answer the question on the validity of this property for the superposed measure $\eta$ mentioned above, and which was left open in [70, Remark 2.10]. Moreover, we deduce the strong Markov property for the nonlinear Markov processes associated with nonlinear Fokker-Planck equation from [6].
@@
    (see Section 2.2) Starting from a solution $(\mu_{t})_{t\geq 0}$ to the linear Fokker-Planck equation (2.3), for each $N>0$ we use the theory of generalized Dirichlet forms for time-dependent operators on $L^{1}$, as developed in [67] (see also [68] and [52]), in order to construct a time-homogenized right process on a subset of the product space $(0,N)\times\mathbb{R}^{d}$, which is of full $\mu_{t}(dx)dt$-measure.
@@ 2.1 The superposition principle and the ”simple” Markov property
Then, according to [35, Theorem 1.1], there exists $\eta\in\mathcal{P}\left(C\left([s,T];\mathbb{R}^{d}\right)\right)$ a solution to the canonical martingale problem associated to $({\sf L}_{t})_{t\in[s,T]}$ on $[s,T]$ such that the one-dimensional time marginals $(\eta_{t})_{t\in[s,T]}$ satisfy $\eta_{t}=\mu_{t},t\in[s,T]$. Moreover, by [70, Proposition 2.8], for each $x\in\mathbb{R}^{d}$ there exists a probability $\eta^{x}\in\mathcal{P}\left(C\left([s,T];\mathbb{R}^{d}\right)\right)$ such that
@@
However, there are relevant examples for which (2.9) is not fulfilled, but for which the relaxed condition (2.8) is valid. Other key results in [70] remain valid if we assume (2.8) instead of (2.9), this being the case of [70, Lemma 2.12] which we shall use below in the proof of 2.7.
@@
The statement follows by [70, Lemma 2.12], choosing $T=N$ and
@@
However, let us point that the proof of 2.8 requires a light adaptation. More precisely, recall first that [70, Remark 2.3] is fundamentally [2, Lemma 8.1.2]. The latter result, remains nevertheless valid if we replace the assumption (2.9) from [70] with the relaxed one (2.8) from [35], with the mention that, under (2.8), the tightness of $(\mu_{t})_{t\in[s,T]}$ required in [2] is ensured by [35, Proposition 2.2 and Proof of Theorem 1.1]; also, note that [35, Proposition 2.2] is valid without the assumption that $(\mu_{t})_{t}$ is weakly continuous, since the latter was used only for applying the Gronwall’s lemma for continuous functions, but the Gronwall’s lemma is valid for merely measurable functions as well, see e.g. [45, Theorem 5.1], and which can be employed in [35, Proposition 2.2].
@@
For the reader convenience let us rapidly prove (2.10), following a standard argument from e.g. [45, Theorem 4.2]: Let $t\in[0,T)$ and $F\in\mathcal{F}_{0}(t)$ with $\eta(F)>0$ be fixed and consider the following probabilities
@@
Then, it follows precisely as in [45, Theorem 4.2] that the probability measures
@@
Note that in [45, Theorem 4.2], the simple Markov property is deduced under the much stronger assumption requiring that the martingale problem has at most one solution for every given initial distribution which even implies that the strong Markov property holds. We stress that above we proved the simple Markov property (2.10) under assumption $\mathbf{H}_{!\mu}$ which is much weaker than the mentioned uniqueness assumption in [45, Theorem 4.2]. However, the strong Markov property can no longer be deduced under this weaker uniqueness assumption by the arguments in [45, Theorem 4.2]. In this paper we are fundamentally interested in proving the strong Markov property as well as other useful regularity properties for the solution the martingale problem associated to the linear Fokker-Planck equation (2.3), not assuming we have uniqueness for the martingale problem under any initial condition. Moreover, we extend these results to the non-linear case as well.
@@ 3. $\mathbf{H^{\sqrt{u}}}$
According to [67, Theorem 1.7, Remark 1.18, Theorem 1.9, Proposition 1.10, Remark 1.11, and Remark 1.12] we have the following collection of fundamental results:
@@ 2.4.1 Some technical lemmas needed in the proofs
This claim follows mostly from [16]. More precisely, by [16, Theorem 2.2 and Remark 2.3], we can directly deduce that a resolvent of sub-Markov kernels $\tilde{\mathcal{U}}$ exists having all the other claimed properties satisfied. Thus, we only need to prove the Markov property pointwise on $0,\infty)\times\mathbb{R}^{d}$. To this end we use 6.53 which is also taken from [16], as follows: Let
@@
Step OP2.1: Firs of all, since the uniqueness assumption $\mathbf{H_{\sf\lesssim u}}$ holds, one can apply [70, Lemma 2.12] to deduce that the family $\left(\eta^{N}\right)_{N>0}$ given by 2.5 is a consistent family (see [62] for the notion), hence by the general theory (see again [62, Theorem 4.1]) admits a unique projective limit $\eta\in\mathcal{P}\left(C([0,\infty);\mathbb{R}^{d})\right)$, with one dimensional marginals $(\mu_{t})_{t\in[0,\infty)}$. Let
@@
is the unique weakly continuous solution to the LFPE (2.3) on $[s,N)$ which consists of probability measures, belongs to $\mathcal{A}_{\lesssim u}^{s,N}$, and whose initial condition (at time $s$) is $\eta_{s}^{\nu_{0}}$, for all $N>s$. But $\left(\eta_{t+s}^{\nu_{0}}\right)_{t\geq 0}$ is also such a solution (see 2.5, in fact [70, Proposition 2.7]), so we must have
@@
    Then by [13, Theorem 1.2.11 and assertion 1 of Remark at page 23] we have that $E^{\dagger,00}\in\mathcal{B}(E^{\dagger,0})$ and $U_{1}^{\dagger,0}(\cdot,E^{\dagger,0}\setminus E^{\dagger,00})=0$ on $E^{\dagger,0}$. Moreover, notice that if we set $E^{\prime}:=E^{\dagger}\setminus\{0\}\times E^{\dagger}(0)$ then $E^{\prime}\in\mathcal{A}(\mathcal{U}^{\dagger})$ (use e.g. 2.12, (ii.4)), $\mathcal{U}^{\dagger,0}|_{E^{\prime}}=\mathcal{U}^{\dagger}|_{E^{\prime}}$, and $\mathcal{U}^{\dagger}$ has no branch points since it is the resolvent of a right process on $E^{\dagger}$; see 6.8 and the remark after it. Thus,
@@
    The next step is to notice that by the monotone class argument [13, Lemma 1.2.10], it is sufficient to fix $f,g\in b{\sf Lip}([0,\infty)\times\mathbb{R}^{d})$ with compact support, where $b{\sf Lip}([0,\infty)\times\mathbb{R}^{d})$ denotes the space of bounded and Lipschitz functions on $0,\infty)\times\mathbb{R}^{d}$, and show that
@@
    In fact, the claim follows by Lions-Magenes interpolation lemma [55, Chapter 3, Theorem 3.1], once we show that
@@
    Also, note that by 2.12, (i), on $L^{p}([0,1]\times\mathbb{R}^{d};\overline{\mu}_{1})$ we can identify $U^{\dagger}_{1}$ with $\overline{U}_{1}^{1}$, where $\overline{U}_{1}^{1}$ is given in 2.10. Consequently, by [67, Theorem 1.7], b) and c), we have that
@@
    whilst $\|u\overline{\sf{L}}_{1}^{1}q\|_{L^{2}((0,1)\times\mathbb{R}^{d};\overline{\mu}_{1})}$ is finite since $\overline{\sf{L}}_{1}^{1}q$ is bounded with compact support by [67, Theorem 1.7, (c)], whilst $u$ is bounded on compact sets by $\mathbf{H^{\sqrt{u}}}$.
@@ 3.2 Backward Kolmogorov equations and the parabolic Dirichlet problem
In order to solve (3.1), we adopt the strategy developed by E.B. Dynkin in [43] in order to treat the so called stochastic Dirichlet problem; see also [60, Section 9]. Crucially, we shall rely on the regularity properties of the balayage operator, as derived in 6.23 and 6.24 in Section 6, which essentially allow us to rigorously employ Doob’s maximal inequality for continuous-time martingales, by ensuring the right-continuity of the trajectories. Let us point out that such a regularity argument seems to have been overlooked in the proof of [60, Theorem 9.2.5] in order to guarantee that, using the notation from [60], the martingale $(N_{t})_{t}$ is right-continuous.
@@
Furthermore, by the strong Markov property and using that $\tau$ is a terminal time ([66, p.65]), hence $\tau=t+\tau\circ\theta(t)$ on $\tau>t$ a.s. for every $t\geq 0$, we have
@@
Note that since $X$ is in fact a Hunt process, it follows from [66, Theorem (47.6) and Exercise (47.7)], that $(\mathcal{F}(t))_{t\geq 0}$ is quasi-left continuous (see [66] for details), hence by e.g. [64, pp. 191-192], no $(\mathcal{F}(t))$-martingale can jump at predictable times. Thus, $\mathbb{P}_{s,x}$-a.s. for $(s,x)\in[(0,T)\times D]\cap E^{\prime}$ we have
@@ 4.2 Trevisan’s conditions
    Note that in [70, Theorem 33], the uniqueness part is apparently claimed in the class of weakly continuous solutions which belong to $L^{\infty}([0,N];L^{r}(\mathbb{R}^{d}))$ and are necessarily probability solutions. However, by inspecting the proof of [70, Theorem 33], one can see that uniqueness holds among all weakly continuous solutions to (2.3) whose densities belongs to $L^{\infty}([0,N];L^{r}(\mathbb{R}^{d}))$.
@@
    We emphasize that the uniqueness is guaranteed in the space $u\in L^{\infty}([0,N];L^{r}(\mathbb{R}^{d}))$, whilst (4.23) is a byproduct of the construction of the solution; see [70, Subsection 3.3] and the discussion on page 20.

### 11 Mar 2026 v2

@@ Abstract
Keywords: Nonlinear Fokker–Planck–Kolmogorov equation; McKean–Vlasov stochastic differential equation; Superposition principle; Right process; Nonlinear Markov process; Porous media equation
Mathematics Subject Classification (2020): 60J45, 60J40, 60J25, 35Q84, 35K55, 60J60; 60J35,
@@ 1 Introduction
where $W$ is a $d$-dimensional standard Brownian motion whilst $\mathcal{L}_{X(t)}$ denotes the law of $X(t)$, $t\geq 0$. The converse implication, namely that the one-dimensional time marginals of a weak solution to the McKean-Vlasov SDE (1.2) solves the (non)linear Fokker-Planck equation (1.1), is just a simple consequence of Ito’s formula. Thus, under mild assumptions on the coefficients, there is a well-understood correspondence between (1.1) and (1.2). The hard implication is, of course, the one that comes from the superposition principle, which was first obtained in the linear case (i.e. the coefficients do not depend on $\mu_{t}$) by [1] and [45], followed up by [64], [35], as well as by other contributions. The extension of the superposition principle to the nonlinear setting and the corresponding McKean-Vlasov SDEs was realized in [3], [4]. We refer to [36] for a comprehensive study of the theory of linear Fokker-Planck-Kolmogorov equations, to [39] for a detailed presentation of McKean-Vlasov SDEs and their applications, and to [6] for the analysis of nonlinear Fokker-Planck equations and their probabilistic counterpart which are nonlinear Markov processes in the sense of McKean; see also the seminal paper [54], as well as the very recent work [59] in which a general recipe is developed how to prove that the family of laws of solutions to a McKean-Vlasov stochastic differential equation with merely measurable coefficients (also only measurable in the measure variable) indexed by a large enough set of initial distributions form a nonlinear Markov process in the sense of McKean. This applies to a large number of examples (see [6] and the references therein), in particular, to also those in [8], [7], where the underlying nonlinear Fokker-Planck equation is the parabolic $p$-Laplace equation and the Leibenson equation respectively, with corresponding non-linear Markov process being the $p$-Brownian motion and the Leibenson process. See also [2] for a gradient flow theory in metric spaces and the superposition principle for the continuity equation.
Another fundamental feature of the superposition principle is that, under a weak uniqueness assumption, the martingale solution denoted above by $\eta$ enjoys the ”simple” Markov property; see Section 2.1, more precisely (2.10). Thus, under very mild assumptions on the coefficients, the correspondence of (non)linear Fokker-Planck equations and McKean Vlasov SDEs could be analyzed through the lens of Markov processes. However, it is well known that the ”simple” Markov property (2.10) is typically not enough in order to benefit from the rich theory of Markov processes, since additional regularity properties of the process, like the strong Markov property, are necessary. This leads us to the formulation of the first main goal which we achieve in this paper:
@@
As indicated above, a right process automatically has the strong Markov property, in particular we answer the question on the validity of this property for the superposed measure $\eta$ mentioned above, and which was left open in [64, Remark 2.10]. Moreover, we deduce the strong Markov property for the nonlinear Markov processes associated with nonlinear Fokker-Planck equation from [6].
@@
    (see Section 2.2) Starting from a solution $(\mu_{t})_{t\geq 0}$ to the linear Fokker-Planck equation (2.3), for each $N>0$ we use the theory of generalized Dirichlet forms for time-dependent operators on $L^{1}$, as developed in [61] (see also [62] and [50], as well as the classical work [47] and [53]), in order to construct a time-homogenized right process on a subset of the product space $(0,N)\times\mathbb{R}^{d}$, which is of full $\mu_{t}(dx)dt$-measure.
@@ 2.1 The superposition principle and the ”simple” Markov property
Then, according to [35, Theorem 1.1], there exists $\eta\in\mathcal{P}\left(C\left([s,T];\mathbb{R}^{d}\right)\right)$ a solution to the canonical martingale problem associated to $({\sf L}_{t})_{t\in[s,T]}$ on $[s,T]$ such that the one-dimensional time marginals $(\eta_{t})_{t\in[s,T]}$ satisfy $\eta_{t}=\mu_{t},t\in[s,T]$. Moreover, by [64, Proposition 2.8], for each $x\in\mathbb{R}^{d}$ there exists a probability $\eta^{x}\in\mathcal{P}\left(C\left([s,T];\mathbb{R}^{d}\right)\right)$ such that
@@
However, there are relevant examples for which (2.9) is not fulfilled, but for which the relaxed condition (2.8) is valid. Other key results in [64] remain valid if we assume (2.8) instead of (2.9), this being the case of [64, Lemma 2.12] which we shall use below in the proof of 2.7.
@@
The statement follows by [64, Lemma 2.12], choosing $T=N$ and
@@
However, let us point that the proof of 2.8 requires a light adaptation. More precisely, recall first that [64, Remark 2.3] is fundamentally [2, Lemma 8.1.2]. The latter result, remains nevertheless valid if we replace the assumption (2.9) from [64] with the relaxed one (2.8) from [35], with the mention that, under (2.8), the tightness of $(\mu_{t})_{t\in[s,T]}$ required in [2] is ensured by [35, Proposition 2.2 and Proof of Theorem 1.1]; also, note that [35, Proposition 2.2] is valid without the assumption that $(\mu_{t})_{t}$ is weakly continuous, since the latter was used only for applying the Gronwall’s lemma for continuous functions, but the Gronwall’s lemma is valid for merely measurable functions as well, see e.g. [44, Theorem 5.1], and which can be employed in [35, Proposition 2.2].
@@
For the reader convenience let us rapidly prove (2.10), following a standard argument from e.g. [44, Theorem 4.2]: Let $t\in[0,T)$ and $F\in\mathcal{F}_{0}(t)$ with $\eta(F)>0$ be fixed and consider the following probabilities
@@
Then, it follows precisely as in [44, Theorem 4.2] that the probability measures
@@
Note that in [44, Theorem 4.2], the simple Markov property is deduced under the much stronger assumption requiring that the martingale problem has at most one solution for every given initial distribution which even implies that the strong Markov property holds. We stress that above we proved the simple Markov property (2.10) under assumption $\mathbf{H}_{!\mu}$ which is much weaker than the mentioned uniqueness assumption in [44, Theorem 4.2]. However, the strong Markov property can no longer be deduced under this weaker uniqueness assumption by the arguments in [44, Theorem 4.2]. In this paper we are fundamentally interested in proving the strong Markov property as well as other useful regularity properties for the solution the martingale problem associated to the linear Fokker-Planck equation (2.3), not assuming we have uniqueness for the martingale problem under any initial condition. Moreover, we extend these results to the non-linear case as well.
@@ 3. $\mathbf{H^{\sqrt{u}}}$
According to [61, Theorem 1.7, Remark 1.18, Theorem 1.9, Proposition 1.10, Remark 1.11, and Remark 1.12] we have the following collection of fundamental results:
@@ 2.4.1 Some technical lemmas needed in the proofs
This claim follows mostly from [17]. More precisely, by [17, Theorem 2.2 and Remark 2.3], we can directly deduce that a resolvent of sub-Markov kernels $\tilde{\mathcal{U}}$ exists having all the other claimed properties satisfied. Thus, we only need to prove the Markov property pointwise on $0,\infty)\times\mathbb{R}^{d}$. To this end we use 6.53 which is also taken from [17], as follows: Let
@@
Step OP2.1: Firs of all, since the uniqueness assumption $\mathbf{H_{\sf\lesssim u}}$ holds, one can apply [64, Lemma 2.12] to deduce that the family $\left(\eta^{N}\right)_{N>0}$ given by 2.5 is a consistent family (see [57] for the notion), hence by the general theory (see again [57, Theorem 4.1]) admits a unique projective limit $\eta\in\mathcal{P}\left(C([0,\infty);\mathbb{R}^{d})\right)$, with one dimensional marginals $(\mu_{t})_{t\in[0,\infty)}$. Let
@@
is the unique weakly continuous solution to the LFPE (2.3) on $[s,N)$ which consists of probability measures, belongs to $\mathcal{A}_{\lesssim u}^{s,N}$, and whose initial condition (at time $s$) is $\eta_{s}^{\nu_{0}}$, for all $N>s$. But $\left(\eta_{t+s}^{\nu_{0}}\right)_{t\geq 0}$ is also such a solution (see 2.5, in fact [64, Proposition 2.7]), so we must have
@@
    Then by [14, Theorem 1.2.11 and assertion 1 of Remark at page 23] we have that $E^{\dagger,00}\in\mathcal{B}(E^{\dagger,0})$ and $U_{1}^{\dagger,0}(\cdot,E^{\dagger,0}\setminus E^{\dagger,00})=0$ on $E^{\dagger,0}$. Moreover, notice that if we set $E^{\prime}:=E^{\dagger}\setminus\{0\}\times E^{\dagger}(0)$ then $E^{\prime}\in\mathcal{A}(\mathcal{U}^{\dagger})$ (use e.g. 2.12, (ii.4)), $\mathcal{U}^{\dagger,0}|_{E^{\prime}}=\mathcal{U}^{\dagger}|_{E^{\prime}}$, and $\mathcal{U}^{\dagger}$ has no branch points since it is the resolvent of a right process on $E^{\dagger}$; see 6.8 and the remark after it. Thus,
@@
    The next step is to notice that by the monotone class argument [14, Lemma 1.2.10], it is sufficient to fix $f,g\in b{\sf Lip}([0,\infty)\times\mathbb{R}^{d})$ with compact support, where $b{\sf Lip}([0,\infty)\times\mathbb{R}^{d})$ denotes the space of bounded and Lipschitz functions on $0,\infty)\times\mathbb{R}^{d}$, and show that
@@
    In fact, the claim follows by Lions-Magenes interpolation lemma [51, Chapter 3, Theorem 3.1], once we show that
@@
    Also, note that by 2.12, (i), on $L^{p}([0,1]\times\mathbb{R}^{d};\overline{\mu}_{1})$ we can identify $U^{\dagger}_{1}$ with $\overline{U}_{1}^{1}$, where $\overline{U}_{1}^{1}$ is given in 2.10. Consequently, by [61, Theorem 1.7], b) and c), we have that
@@
    whilst $\|u\overline{\sf{L}}_{1}^{1}q\|_{L^{2}((0,1)\times\mathbb{R}^{d};\overline{\mu}_{1})}$ is finite since $\overline{\sf{L}}_{1}^{1}q$ is bounded with compact support by [61, Theorem 1.7, (c)], whilst $u$ is bounded on compact sets by $\mathbf{H^{\sqrt{u}}}$.
@@ 3.2 Backward Kolmogorov equations and the parabolic Dirichlet problem
In order to solve (3.1), we adopt the strategy developed by E.B. Dynkin in [42] in order to treat the so called stochastic Dirichlet problem; see also [56, Section 9], or [10]. Crucially, we shall rely on the regularity properties of the balayage operator, as derived in 6.23 and 6.24 in Section 6, which essentially allow us to rigorously employ Doob’s maximal inequality for continuous-time martingales, by ensuring the right-continuity of the trajectories. Let us point out that such a regularity argument seems to have been overlooked in the proof of [56, Theorem 9.2.5] in order to guarantee that, using the notation from [56], the martingale $(N_{t})_{t}$ is right-continuous.
@@
Furthermore, by the strong Markov property and using that $\tau$ is a terminal time ([60, p.65]), hence $\tau=t+\tau\circ\theta(t)$ on $\tau>t$ a.s. for every $t\geq 0$, we have
@@
Note that since $X$ is in fact a Hunt process, it follows from [60, Theorem (47.6) and Exercise (47.7)], that $(\mathcal{F}(t))_{t\geq 0}$ is quasi-left continuous (see [60] for details), hence by e.g. [58, pp. 191-192], no $(\mathcal{F}(t))$-martingale can jump at predictable times. Thus, $\mathbb{P}_{s,x}$-a.s. for $(s,x)\in[(0,T)\times D]\cap E^{\prime}$ we have
@@ 4.1 The square root $H^{1}$-regularity for Fokker-Planck equations
The following regularity result is the first main step towards the main result of this subsection, namely 4.4 from below. For its proof we got inspired from [34], where the uniform ellipticity is employed to prove the regularity of invariant measures for time-independent operators.
@@ 4.2 Trevisan’s conditions
    Note that in [64, Theorem 33], the uniqueness part is apparently claimed in the class of weakly continuous solutions which belong to $L^{\infty}([0,N];L^{r}(\mathbb{R}^{d}))$ and are necessarily probability solutions. However, by inspecting the proof of [64, Theorem 33], one can see that uniqueness holds among all weakly continuous solutions to (2.3) whose densities belongs to $L^{\infty}([0,N];L^{r}(\mathbb{R}^{d}))$.
@@
    We emphasize that the uniqueness is guaranteed in the space $u\in L^{\infty}([0,N];L^{r}(\mathbb{R}^{d}))$, whilst (4.23) is a byproduct of the construction of the solution; see [64, Subsection 3.3] and the discussion on page 20.